# OmniHands: Towards Robust 4D Hand Mesh Recovery via A Versatile Transformer

### v3

@@ 4.5 Comparison with state-of-the-art methods
Quantitive Comparison on In-the-wild Datasets. We report the results on three in-the-wild datasets: Arctic [13], HIC [55] and RenderIH [31], to compare the performance at in-the-wild scenarios. Arctic contains images of two hand interacting with objects, HIC is a small dataset of two hand interaction sequences and RenderIH contains images of complicated two hand interactions under various light conditions. We only use samples in which both hands are visible. For fair comparison, temporal data or multi-view data are not used in these experiments, a single image is taken as input for each case. In Table 4 we compare the results with IntagHand [32], InterWild [38] and DIR [47]. IntagHand [32] and DIR [47] are two open-source inter-hand reconstruction models that have demonstrated excellent performance in training and testing on Interhand2.6m [39]. InterWild [38] is a SOTA method specifically designed for two-hand recovery in the wild. To demonstrate the efficiency of using large training data, we report results of models trained under two training set. In all three unseen datasets, our method achieves the best performance and leads in all metrics. It can also be found that by using large training set, our model can handle situations in more diverse scenarios, resulting in higher accuracy in all three testing sets. HIC [55] includes interactive gestures that are not present in Interhand2.6m. However, the scenes are relatively simple and the test samples are few, allowing other methods to also achieve good results. In HIC our method improves 16.33% (8.66 mm $vs.$ 10.35 mm) on Ra-MPJPE and 29.47% (4.81 mm $vs.$ 3.34 mm) on Pa-MPJPE. Arctic [13] includes a large number of images of two hands interacting with objects, which significantly differ from the data in Interhand2.6m. Additionally, there are many instances where parts of the hands are occluded, making the situation more complex. In this dataset, other methods performed poorly, whereas our method maintained robustness and achieved excellent results, improving 45.46% (13.57 mm $vs.$ 24.88 mm) on Ra-MPJPE and 33.61% (7.98 mm $vs.$ 12.02 mm) on Pa-MPJPE. In RenderIH [31], the diversity of lighting and backgrounds places high demands on the cross-scene robustness of the model. Our proposed Omnihands obtains 13.89% (13.71 mm $vs.$ 15.92 mm) improvement on Ra-MPJPE and 13.00% (9.17 mm $vs.$ 10.54 mm) on Pa-MPJPE.
Figure 5: Qualitive results of OmniHands with multi-view inputs on Interhand2.6m [39] and Arctic [13]. In each row, we present pairs of input and output images from 4 different camera views. The images on the left side of each pair serve as the multi-view sequence input.
@@
Comparison on Single hand Multi-view Dataset. To demonstrate the performance of our model when taking multi-view images as input, in Table 3 we compare with other multi-view single hand methods on Dex-YCB [4]. We also compare with Multi-view pose estimation reported in [63]. It’s worth mentioning that other methods require camera parameters as input, which makes it easier for them to align different views but loses flexibility. Our method, on the other hand, does not need camera parameters as input. It only requires multiple images from different viewpoints, eliminating the tedious camera calibration steps while achieving higher accuracy than other methods.
4.6 Ablation Study
@@
We evaluate the importance of each module in Omnihands by albation experiments in temporal input conditions, using temporal dataset Interhand2.6m 30fps as training set and tesing set. Moreover, we also conduct evaluations on the impact of camera numbers when taking multi-view images as input, using Interhand2.6m 5fps multi-view dataset.
4.6.1 Module Ablation
@@
ViT Baseline. We trained a baseline using ViT and a transformer decoder on full training set. Position map embedding, cross-hand feature interaction and temproal/multi-view interaction are not employed on the baseline, so the Bimanual Decoder, Collective Decoder and Individual Encoder are removed. We show the results in Table 6. Although ViT has high performance in detecting single hand, the occlusion caused by hands interaction reduces its predictive ability. Besides, without cross-hand knowledge, it separately predict positions of two hand, causing huge errors in relative position.
Figure 6: Ablation study of RAT. Two views of predicted hand meshes are visualized to demonstrate details of relative position. ’Full Model’ is our model with all modules. ’w/o RAT’ means that the position embedding is removed.
@@
Figure 7: Ablation study on occlusion case. ’Full Model’ is our model with all modules. ’w/o Temporal Fusion’ means that the temporal integrating module is removed.
@@
Effectiveness of Relation-aware Tokenization. By removing position token map embedding in RAT(+Cross S w/o RAT), the model lost global position information during inference, which means that the model could only infer two-hand relative position from visual information. In InterHand2.6m [39], the two hands are relatively close and the 2D scales of each hand are similar. Therefore, even without a global position information, the approximate positions of both hands can be inferred from the input image. We test cases when two hands are relatively far away and cases when 2D scales of each hand are different. As shown in Table 6, without position token map, the model performs poor in predicting two-hand relative position.
@@
4.6.2 Camera view Ablation
In multi-view reconstruction, the number of input camera views is important to the resulting accuracy. We tried to figure out the impact of camera views, and the experimental results are shown in Table 7. We divided the experimental results into two groups based on whether mixed training data was used. ’I’ indicates the use of only Interhand2.6m[39] as the training data, while ’I+R+A’ represents the use of a mixed training dataset comprising Interhand2.6m, Re:Interhand[41], and Arctic[13].
@@ 4.6.2 Camera view Ablation
Table 6: Ablation studies of the RAT and FIR modules on InterHand2.6M.
@@
Table 7: Results on Interhand2.6m using different number of camera views. ”I” means only using Interhand2.6m training set, and ”I+R+A” means using Interhand2.6m, Re:Interhand and Arctic as training set.
@@
5 Limitaions
Our method utilizes ViT-h as the backbone, which comprises 32 transformer layers, resulting in higher computational demands compared to other CNN-based methods. During inference, running on a single A800 GPU with Python code, our model achieves 16.05 FPS, indicating that it is not suitable for real-time rendering. Additionally, because we use external detection models to locate hands in images and regress the bounding boxes, the inference speed is also constrained by the performance of the detection models. We will focus on accelerating our model by reimplementing it in C++ instead of Python and optimizing the computation process to enable real-time rendering.
6 Conclusions
In this paper, we presented OmniHands, a robust solution for recovering interactive hand meshes and their relative movements from monocular inputs. OmniHands is a transformer-based method with novel tokenization and feature fusion strategies. To this end, we introduced the Relation-aware Two-Hand Tokenization (RAT) method, which embeds positional relation information into hand tokens. This allows our network to handle both single-hand and two-hand inputs, leveraging their relative positions to reconstruct detailed hand interactions in real-world scenarios. Moreover, this tokenization approach enhances feature fusion by indicating the relative positions of the hands. To further this goal, we developed a 4D Interaction Reasoning (FIR) module that fuses hand tokens in 4D with attention, decoding them into 3D hand meshes and their temporal movements. The effectiveness of OmniHands is demonstrated through extensive evaluations on benchmark datasets, with results from in-the-wild videos and real-world scenarios showcasing its superior performance in interactive hand reconstruction.

### v4

@@ 4.5. Comparison with state-of-the-art methods
In all three unseen datasets, our method achieves the best performance and leads in all metrics. It can also be found that by using large training set, our model can handle situations in more diverse scenarios, resulting in higher accuracy in all three testing sets. HIC (Tzionas et al., 2016) includes interactive gestures that are not present in Interhand2.6m. However, the scenes are relatively simple and the test samples are few, allowing other methods to also achieve good results. In HIC our method improves 16.33% (8.66 mm $vs.$ 10.35 mm) on Ra-MPJPE and 29.47% (4.81 mm $vs.$ 3.34 mm) on Pa-MPJPE. Arctic (Fan et al., 2023) includes a large number of images of two hands interacting with objects, which significantly differ from the data in Interhand2.6m. Additionally, there are many instances where parts of the hands are occluded, making the situation more complex. In this dataset, other methods performed poorly, whereas our method maintained robustness and achieved excellent results, improving 45.46% (13.57 mm $vs.$ 24.88 mm) on Ra-MPJPE and 33.61% (7.98 mm $vs.$ 12.02 mm) on Pa-MPJPE. In RenderIH (Li et al., 2023), the diversity of lighting and backgrounds places high demands on the cross-scene robustness of the model. Our proposed OmniHands obtains 13.89% (13.71 mm $vs.$ 15.92 mm) improvement on Ra-MPJPE and 13.00% (9.17 mm $vs.$ 10.54 mm) on Pa-MPJPE.
@@
Comparison on Single hand Multi-view Dataset. To demonstrate the performance of our model when taking multi-view images as input, in Table 3 we compare with other multi-view single hand methods on Dex-YCB (Chao et al., 2021). For a fair comparison, our model is trained on Dex-YCB. We also compare with Multi-view pose estimation reported in (yang2023poem). It’s worth mentioning that other methods require camera parameters as input, which makes it easier for them to align different views but loses flexibility. Our method, on the other hand, does not need camera parameters as input. It only requires multiple images from different viewpoints, eliminating the tedious camera calibration steps while achieving higher accuracy than other methods.
4.6. Ablation Study
@@
We evaluate the importance of each module in OmniHands by ablation experiments in temporal input conditions, using temporal dataset Interhand2.6m 30fps as training set and testing set. Moreover, we also conduct evaluations on the impact of camera numbers when taking multi-view images as input, using Interhand2.6m 5fps multi-view dataset.
4.6.1. Module Ablation
@@
ViT Baseline. We trained a baseline using ViT and a transformer decoder on the full training set. Position map embedding, cross-hand feature interaction and temporal/multi-view interaction are not employed on the baseline, so the Bi-manual Decoder, Collective Decoder, and Individual Encoder are removed. We show the results in Table 9. Although ViT has high performance in detecting single hand, the occlusion caused by hands interaction reduces its predictive ability. Besides, without cross-hand knowledge, it separately predicts the positions of two hands, causing huge errors in relative position.
@@
Figure 6. Ablation study on occlusion case. ’Full Model’ is our model with all modules. ’w/o Temporal Fusion’ means that the temporal integrating module is removed.
A figure.
@@
Effectiveness of Relation-aware Tokenization. By removing position token map embedding in RAT(+Cross S w/o RAT), the model lost global position information during inference, which means that the model could only infer two-hand relative positions from visual information. In InterHand2.6m (Moon_2020_ECCV_InterHand2.6M), the two hands are relatively close and the 2D scales of each hand are similar. Therefore, even without a global position information, the approximate positions of both hands can be inferred from the input image. We test cases when two hands are relatively far away and cases when 2D scales of each hand are different. As shown in Table  7, without position token map, the model performs poor in predicting two-hand relative position.
Figure 7. Ablation study of RAT. Two views of predicted hand meshes are visualized to demonstrate details of relative position. ’Full Model’ is our model with all modules. ’w/o RAT’ means that the position embedding is removed.
A figure.
@@
4.6.2. Ablation on training data.
In this section, we conduct ablation studies to evaluate the impact of scaling up training data. We benchmark our method on three datasets: InterHand2.6m (Moon_2020_ECCV_InterHand2.6M), HIC (Tzionas et al., 2016), and Arctic (Fan et al., 2023), and compare against two SOTA baselines: IntagHand (Li et al., 2022) and InterWild (Moon, 2023). The results are presented in Table 11. For each method, we train the models on both the InterHand2.6M 5fps dataset and mixed datasets. Specifically, the mixed datasets for OmniHands include InterHand2.6M, Re:InterHand, and Static Single Hand Datasets. The static single hand datasets simulate a pair of widely separated hands by randomly combining two single-hand images. However, such input is not suitable for IntagHand or InterWild, as their input requires images that contain both hands. Therefore, for IntagHand and InterWild, the mixed datasets comprise InterHand2.6M and Re:InterHand. Concerning the mixed datasets, while OmniHands leverages additional single-hand datasets during training, all methods are trained on the same interacting-hand datasets. As the prediction of relative hand positions is learned from the interacting-hand data, the MRRPE metric serves as a valid indicator for comparing the interacting-hand reconstruction performance among the methods.
The experimental results demonstrate that OmniHands benefits from the increased data diversity, exhibiting improved performance when trained with the hybrid dataset. In contrast, IntagHand suffers performance degradation with additional training data. This is likely due to its architectural limitations, which prevent it from effectively generalizing under increased data scale and complexity, such as varying scenes and lighting conditions, ultimately undermining its reliance on strong priors. Overall, these results validate that OmniHands consistently outperforms SOTA methods, regardless of whether the model is trained on a single-domain dataset or a more diverse hybrid dataset.
4.6.3. Camera view Ablation
In multi-view reconstruction, the number of input camera views is important to the resulting accuracy. We tried to figure out the impact of camera views, and the experimental results are shown in Table 10. We divided the experimental results into two groups based on whether mixed training data was used. ’I’ indicates the use of only Interhand2.6m(Moon_2020_ECCV_InterHand2.6M) as the training data, while ’I+R+A’ represents the use of a mixed training dataset comprising Interhand2.6m, Re:Interhand(Moon et al., 2024), and Arctic(Fan et al., 2023).
@@ 4.6.3. Camera view Ablation
Table 9. Ablation studies of the RAT and FIR modules on InterHand2.6M.
@@
Table 10. Results on Interhand2.6m using different number of camera views. ”I” means only using Interhand2.6m training set, and ”I+R+A” means using Interhand2.6m, Re:Interhand and Arctic as training set.
@@
Table 11. Albation study on training dataset. Models denoted with ’-i’ is trained on Interhand2.6m. Models denoted with ’-m’ is trained on mixed datasets.
Dataset Methods MPJPE MPVPE MRRPE
Inter2.6m 5fps IntagHand-i 8.79 9.02 27.74
IntagHand-m 10.55 10.98 30.04
InterWild-i 14.76 13.24 26.25
InterWild-m 14.45 13.01 25.68
Ours-i 7.49 7.72 24.58
Ours-m 6.77 6.96 22.01
HIC IntagHand-i 12.06 12.77 36.46
IntagHand-m 12.54 13.32 35.41
InterWild-i 11.23 11.65 30.35
InterWild-m 11.13 11.54 30.59
Ours-i 9.47 10.35 26.84
Ours-m 8.66 9.31 26.56
Arctic IntagHand-i 24.88 26.30 203.14
IntagHand-m 27.64 32.78 225.49
InterWild-i 22.74 23.92 143.19
InterWild-m 21.85 22.98 134.69
Ours-i 14.21 15.12 112.81
Ours-m 13.57 14.51 105.43
5. Limitaions
Although OmniHands has demonstrated high accuracy in offline video-based reconstruction, there remain several challenges that currently hinder its deployment for real-time interacting-hand reconstruction. Our method utilizes ViT-h as the backbone, resulting in higher computational demands compared to CNN-based methods, which is a common issue with transformer-based models, as seen in HaMeR (Pavlakos et al., 2023) and Hamba (dong2024hamba). During inference, running on a single A800 GPU with Python code, our model achieves 16.05 FPS, indicating that it is not suitable for real-time rendering, which requires an efficiency of at least 25 to 30 FPS. Additionally, since our approach relies on external detection models to localize hands and regress bounding boxes, the overall inference speed is also limited by the performance of these detection models. Another challenge in real-time reconstruction arises from the fact that we predict the middle frame, which introduces a minimum delay of 4 frames in a live video stream, as the model requires access to 4 future frames as input. We will focus on accelerating our model by reimplementing it in C++ instead of Python and optimizing the computation process to enable real-time rendering.
6. Conclusions
In this paper, we presented OmniHands, a robust solution for recovering interactive hand motions and their relative movements from monocular inputs. OmniHands is a transformer-based method with novel tokenization and feature fusion strategies. To this end, we introduced the Relation-aware Two-Hand Tokenization (RAT) method, which embeds positional relation information into hand tokens. This allows our network to handle both single-hand and two-hand inputs, leveraging their relative positions to reconstruct detailed hand interactions in real-world scenarios. Moreover, this tokenization approach enhances feature fusion by indicating the relative positions of the hands. To further this goal, we developed a 4D Interaction Reasoning (FIR) module that fuses hand tokens in 4D with attention, decoding them into 3D hand meshes and their temporal movements. The effectiveness of OmniHands is demonstrated through extensive evaluations on benchmark datasets, with results from in-the-wild videos and real-world scenarios showcasing its superior performance in interactive hand reconstruction.